OpenStudy (anonymous):

Verify the identity (1 + tan^2u)(1 - sin^2u) =

OpenStudy (anonymous):

Seriously need help please I need a thorough explanation of how this will look cause I'm extremely lost :(

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@dan815

OpenStudy (bibby):

use your trig identies sin^2x+cos^2x=1 tan = sin/cos

OpenStudy (anonymous):

can you walk me through it

OpenStudy (bibby):

no my son ned me 2 taqke him to the school...

OpenStudy (anonymous):

ohh ok :/

OpenStudy (bibby):

just an fyi, u didnt even include the right hand side of the identity

OpenStudy (anonymous):

that's all I was given

OpenStudy (amorfide):

it is not an identity as there is nothing on the right hand side to verify the left hand side

OpenStudy (bibby):

oh, I figured we were proving an equality. again this is just identity manipulation. it's algebra

OpenStudy (anonymous):

@bibby @amorfide sorry its actually (1 + tan2u)(1 - sin2u) = 1

OpenStudy (amorfide):

\[(1+\tan^{2}u)(1-\sin^{2}u)=1\]

OpenStudy (anonymous):

and the 2s are powers

OpenStudy (amorfide):

is this your question?

OpenStudy (bibby):

http://puu.sh/gSMnF.png

OpenStudy (anonymous):

yes amor

OpenStudy (amorfide):

well you know that \[\tan(u)=\frac{ \sin(u) }{ \cos(u) }\] from \[\sin^{2}(u)+\cos^{2}(u)=1\] we know that \[1-\sin^{2}(u)=\cos^{2}(u)\]

OpenStudy (amorfide):

so we can work out \[\tan^{2}(u)=\frac{ \sin^{2}(y) }{ \cos^{2}(u) }\]

OpenStudy (amorfide):

so replace tan squared replace 1-sin squared

OpenStudy (amorfide):

multiply out of the brackets

OpenStudy (amorfide):

and you can work it out from that I am sure

OpenStudy (amorfide):

let me know if you still need help after doing this

OpenStudy (anonymous):

so the explanation so far is tan^2(u)= sin^2(u)/cos^2(u) which changes the equation to (1+sin^2(u)/co^2(u))*(cos^2(u)=1

OpenStudy (amorfide):

yes

OpenStudy (amorfide):

now multiply out of the brackets and you can simplify

OpenStudy (anonymous):

and that's now verified?

OpenStudy (amorfide):

you will end up with \[\cos^{2}(u)+\sin^{2}(u)=1\] which is true since that is also an identity

OpenStudy (amorfide):

that is now verified since left hand side is the same as the right hand side

OpenStudy (amorfide):

let me know if you do not understand something

OpenStudy (anonymous):

isn't that the same equation as you get to use for 1-sin^2(u)

OpenStudy (amorfide):

you use the trig identity for \[\sin^{2}z+\cos^{2}x=1\] to work out equivalent expressions for the ones in your question, and it just happens to come out as the trig identity we all know and love it does not always work out that way though

OpenStudy (anonymous):

so can you explain how you broke down (1+sin^2(u)/cos^2(u)) * (cos^2(u) into sin^2(u)+cos^2(u)=1

OpenStudy (amorfide):

that should be a plus my bad

OpenStudy (anonymous):

im going to take a smoke so ill be back in 2 minutes or so if you don't mind

OpenStudy (amorfide):

\[(1+\frac{ \sin^{2}(u) }{ \cos^{2}(u) })(\cos^{2}(u))\]

OpenStudy (amorfide):

\[\cos^{2}(u) + \frac{ \sin^{2}(u)\cos^{2}(u) }{ \cos^{2}(u) }\]

OpenStudy (amorfide):

\[\frac{ \cos^{2}(u) }{ \cos^{2}(u) }=1\]

OpenStudy (amorfide):

\[\cos^{2}(u)+\sin^{2}(u)\]

OpenStudy (amorfide):

so this is our left hand side, and we know the right hand side is 1

OpenStudy (amorfide):

using the trig identity \[\cos^{2}(u)+\sin^{2}(u)=1\] we have 1=1

OpenStudy (anonymous):

can you explain your transition from cos^2(u)+sin^2(u) cos^2(u)/cos 2 (u) to cos^2(u)/cos^2(u)=1

OpenStudy (amorfide):

\[\cos^{2}(u)+ \frac{\sin^{2}(u) \cos^{2}(u) }{ \cos^{2}(u) }\] this is what i had

OpenStudy (amorfide):

there was no transition

OpenStudy (amorfide):

i was showing that the fraction cancels out to just sin squared

OpenStudy (amorfide):

because anything divided by itself is 1 so it was the same as sin^2(u) times 1

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

@amorfide one last thing what happened to the 1+

OpenStudy (amorfide):

what do you mean

OpenStudy (amorfide):

we multiplied by cos squared remember

OpenStudy (anonymous):

(1+sin^2(u)/cos^2(u)) where did the +1 in this go

OpenStudy (anonymous):

@amorfide

OpenStudy (amorfide):

you missed out the other half of your trig you had two brackets to start with (1+tan^2(u))(1-sin^2(u)) 1+tan^2(u) = 1+ sin^2(u)/cos^2(u) 1-sin^2(u)= cos^2(u) we multiply them together (1+sin^2(u)/cos^2(u)) x (cos^2(u)) multiply everything in the first bracket by cos^2(u)

OpenStudy (anonymous):

ohh I tryied to do that wrong I get it now I was tryin to combine the two sides of the equation cause im stupid

OpenStudy (anonymous):

(1+tan^2 u)(1-sin^2 u)= (sec^2u)(cos^2)= (1/cos^2)(cos^2u)=1

OpenStudy (anonymous):

1+tan^2u = sec^2 u 1-sin^2= cos^2 sec^2=1/cos^2